sowmya takes 6 days less than the no of days taken by Bhagya to complete a piece of work if both sowmya and bhagya together can complete the same work in 4 days in how many days wil bhagya alone complete the work
Answer by mananth(16949)

sowmya takes 6 days less than the no of days taken by Bhagya to complete a piece of work if both sowmya and bhagya together can complete the same work in 4 days in how many days wil bhagya alone complete the work
..
let bhagya take x days
sowmya takes x-6 days
...
bhagya does 1/x of the job in 1 fday
sowmya does 1/(x-6) of the job in 1 day.
..
together they can do 1/x + 1/(x-6) =f the job in 1 day
LCM = x(x-6)
(x-6+x)/x(x-6)
2x-6/x(x-6)
so they takes x(x-6)/2x-6 hours
x^2-6x/(2x-6)=4
x^2-6x=4(2x-6)
x^2-6x=8x-24
x^2-14x+24=0
x^2-12x-2x+24=0
x(x-12)-2(x-12)=0
(x-12)(x-2)=0
x= 12 days taken by bhagya alone
m.ananth@hotmail.ca
..